# LeLaR: The First In-Orbit Demonstration of an AI-Based Satellite Attitude Controller

### v2

@@ Abstract
In-Orbit Demonstration , Artificial Intelligence , Deep Reinforcement Learning , Nanosatellites , CubeSats , Attitude Control
@@ 1 Introduction and Background
In recent years, Machine Learning (ML) and Artificial Intelligence (AI)-based methods have increasingly entered engineering disciplines, particularly robotics, with notable progress in quadrupedal [DBLP:journals/corr/abs-2109-11978] and humanoid robots such as Agile Justin developed at the German Aerospace Center (DLR) [R_stel_2025]. This influence is now expanding into the aerospace sector, where AI-driven methods are being explored for highly nonlinear applications, including robust attitude control for launch vehicles [aerospace12030203, Xue2023], combustion chamber regulation [Hoerger2025], UAV control [aerospace12080684, Mohiuddin2025], and satellite attitude control [doi:10.2514/6.2025-1145, upm90311, retagne2024adaptive].
The motivation stems from the increasing autonomy and complexity required by modern spacecraft, which demand control strategies capable of adapting to dynamic environments while minimizing human intervention. Deep Reinforcement Learning (DRL), a variant of classical Reinforcement Learning [sutton1998reinforcement], has shown considerable promise in this context. For example, for launch vehicles, DRL-based controllers using the Soft Actor–Critic [haarnoja2018soft] algorithm have been applied in simulation to achieve load-relief attitude control under aerodynamic disturbances [aerospace12030203]. Similarly, DRL-based adaptive controllers for quad-tiltrotor UAVs improve disturbance rejection and tracking performance compared to classical methods [aerospace12080684] in simulation.
For satellites, DRL has been used in simulation to maintain operational capability under unknown ADCS failures, where a Deep Deterministic Policy Gradient (DDPG) [lillicrap2015continuous]-based controller reorients solar panels toward the Sun without requiring explicit fault models [doi:10.2514/6.2025-1145]. Other work demonstrates the feasibility of real-time DRL-based magnetic attitude control on embedded hardware on ground [upm90311]. In addition, Proximal Policy Optimization (PPO) [schulman2017proximal]-based controllers using stacked observations, similar to our observation space design from [djebko2023learning], achieve robust attitude control across varying satellite masses in simulation, outperforming classical PID control in settling time and stability [retagne2024adaptive]. The Hybrid Online Policy Adaptation Strategy (HOPAS) project attempted an early in-orbit demonstration of a Reinforcement Learning-based attitude controller on the ESA OPS-SAT CubeSat, with the AI agent designed to modify the PID control signals as a residual command, but the full AI closed-loop control was ultimately restricted to simulation and ground-testing [CarrilloBarrenechea2023, Airbus2022HOPAS]. These developments illustrate the growing interest in DRL and AI-based control methods for aerospace systems characterized by nonlinear dynamics and uncertainties.
@@
Building on our previous works [djebko2023learning, gerlich2023its, djebko2025], we developed an AI-based attitude controller for the InnoCube 3U nanosatellite [aerospace8050127, montenegro2022innocube, baumann2025innocube], developed by the Julius-Maximilians-Universität Würzburg and the Technische Universität Berlin, which successfully launched on January 14, 2025. The AI agent autonomously interacts with a simulated environment to learn a control strategy that maximizes a given reward function. The agent was first trained on a basic control task to obtain a base-agent, which was subsequently post-trained to produce the final flight-agent. The controller was trained entirely on ground, uploaded to InnoCube, and tested in orbit. Since the present work was done as part of the “In-Orbit Demonstrator Learning Attitude Control” (German: In-Orbit-Demonstrator Lernende Lageregelung; LeLaR) project, we also refer to the AI agent as LeLaR. The InnoCube project is funded by the Federal Ministry for Economic Affairs and Energy (BMWE) under grant number (FKZ) 50RU2000.
A disadvantage of DRL lies in the tendency of neural networks to achieve good results in interpolation tasks but failure in extrapolation problems [tsimenidis2020limitationsdeepneuralnetworks]. In that sense, the process of transferring an AI agent trained in simulation to the real target system is called Sim2Real transfer and poses a significant challenge when applying DRL to real-world tasks. For practical applications, it is necessary to vary the boundary conditions during training, e.g., via domain randomization, to avoid so-called Out-of-Distribution errors. This includes, in particular, the moments of inertia, which we have already successfully considered in [djebko2023learning, gerlich2023its] and utilize in this work.
@@ 2 Methodology
As base-agent, we employ the post-trained agent from our prior work [Djebko2025_SubnetworkPolicy], using a split policy network, consisting of a reaction wheel and a magnetorquer subnetwork. It was originally designed to avoid prolonged RW operation in the interval of [350,350][-350,350] rpm to avoid the inaccuracies stemming from the RWs rate estimator when operated in this range. Initial tests in-orbit however revealed that the RW subnetwork alone performed sufficient well, as we will discuss in Section 3. Along with changed safety requirements introduced after deployment but prior to the first test, this observation prompted us to use the RW subnetwork of the agent from [Djebko2025_SubnetworkPolicy] as base and to fine tune it through post-training for standalone use. We refer to the post-trained agent as LeLaR flight-agent or flight-agent in short.
Since [Djebko2025_SubnetworkPolicy] is available only in German, we adapt and summarize the essential elements required to understand the current work and expand on important aspects. This includes the base-agent design and the sensor calibration methodology adapted from our prior work, covering both the physical onboard sensors and the simulation environment used for training of all agents. It is to be noted, that sensor calibration is not a requirement introduced by the DRL procedure but should generally be performed for any satellite mission. It is included here for completeness, as calibration affects both the satellite’s sensors and the quality of the simulation. We then highlight the modifications to the model and reward function, as well as the updated operational requirements and boundary conditions, applied to obtain the flight-agent. Finally, the Safety Cage architecture is described, encompassing the AI agent to ensure safe and reliable operation in case of a fault.
@@ 2.1 InnoCube Satellite
InnoCube is a 3U CubeSat mission designed as a technology demonstrator for several innovative technologies and has successfully been launched on January 14, 2025 into a 520km sun-synchronous orbit. InnoCube carries two main experiments: “Skip The Harness” (SKITH), a fully wireless data bus for intra-satellite communication and “Fibre-Reinforced Spacecraft Wall for Storing Energy” (Wall#E), an experimental carbon fiber reinforced plastic solid-state battery. Secondary payloads include amateur radio payloads and the “Experiment for Precise Orbit Determination” (EPISODE) experiment, a software defined GNSS receiver. InnoCube is a collaborative effort between the Julius-Maximilians-Universität Würzburg and the Technische Universität Berlin [baumann2025innocube]. The InnoCube satellite is a 3U CubeSat with a mass of 4.2 kg. An overview of the spacecraft’s subsystems is provided in Figure 1.
@@
On-board data handling is distributed across the satellite’s individual hardware nodes. The system is based on the Silicon Labs EFR32FG12 microcontroller, which features an ARM Cortex-M4 processor [efr32fg12]. It integrates a 2.4 GHz RF front end that serves as the foundation for the wireless bus system Skip-the-Harness (SKITH). All spacecraft modules share this wireless data interface. Mission-critical bus functions are implemented within the PCDU, the COM node and the OBC node. These include power management, telemetry and telecommand functionality, as well as general bus operations.
@@ 2.1.1 InnoCube ADCS
As previously mentioned in Section 2.1, all spacecraft modules share the same EFR32FG12 microcontroller. On the ADCS mainboard, the basic circuitry of the microcontroller is extended to accommodate the electrical components needed for the ADCS. These include the gyroscope, the magnetometer, controllers and electrical interfaces for the magentorquers, electrical interfaces to the reaction wheels and other nescessary electrical systems. The used gyroscope to measure the body rates of the satellite is a STMicroelectronics ASM330LHH micro-electromechanical systems (MEMS) gyro [asm330gyro]. The used magnetometer to measure the magnetic vector with respect to the body frame is a three-axis PNI RM3100 magneto-inductive sensor [rm3100mag]. The external sun sensors are not used for the current LeLaR experiments, since we currently only determine the relative attitude using a quaternion integrator based on the measured gyro body rates.
@@ 2.2 Simulation Model
The simulated orbit is based on InnoCube TLE data from CelesTrak [celestrak2025innocube], with variations applied during training. Table 3 summarizes nominal values and offsets.
@@ 2.4.1 Sensor Noise Model
Bias std $\sigma_{b}$ (rad/s) $2.008\times 10^{-2}$ $1.419\times 10^{-2}$ $9.774\times 10^{-2}$
@@ 2.4.3 Residual Dipole Moment Estimation
Residual dipole $\mu$ and gyro bias $b$ were estimated from telemetry by solving a calibration problem based on Euler’s rotational dynamics [BUSCH201573]:
@@
where $\omega_{corrected}=\omega_{measured}+b$. The method was validated using simulated data with known ground truth and then applied to telemetry without actuator influence. Leveraging PyTorch [paszke2019pytorchimperativestylehighperformance] autograd gradient descent and the AdamW optimizer, we obtained:
@@ 2.5 Base-Agent Training
Using our AI agent from [Djebko2025_SubnetworkPolicy] as a base, we derive the flight-agent through post-training, which is conceptually related to transfer learning. The base-agent is initially trained to solve the general attitude control problem. Subsequently, it undergoes post-training with modified boundary conditions and a revised optimization target to produce the flight-agent. This approach leverages the well-initialized policy that already captures the fundamental dynamics of the system, avoiding the need to re-learn this knowledge as compared to training the flight-agent from scratch. As outlined in Section 1 the flight-agent is based off of the RW subnetwork of the base-agent. In the following section, the base-agent training is described. The modifications performed to obtain the flight-agent are explained in Section 2.6.
@@ 2.5.1 Base-Agent Reward Functions
The base-agent was trained with a PPO derivative, SkipPPO [Djebko2025_SubnetworkPolicy], using separate subnetworks for RWs and MTs. The RWs learn to quickly attain the goal attitude, while the MTs learn slow RW momentum management to bring the RW speeds to $\pm$500 rpm (corresponding to $\pm$350 rpm and a safety buffer). The RW reward function of the base-agent rewards quick reduction of attitude error, reduction of angular rates, and stable convergence. The MT reward encourages minimizing RW speed deviations. A combined reward was used during post-training of the base-agent to balance both effects. The RW reward function is shown in Equations (11)–(16).
@@ 2.7.1 Safety Cage
In order to assure that mission critical states resulting from unwanted controller behaviour are not reached during attitude control maneuvers, a Safety Cage architecture as proposed in [aiecsshandbook] is implemented. This construct, recommended by the ECSS Machine Learning Handbook, uses a set of rules, to prevent the system from entering potentially dangerous states. This can be achieved through a set of safety interventions, which will be activated once a certain rule is triggered. Potentially critical failure modes for the InnoCube satellite have been identified using the proposed Failure Mode, Effects and Criticality analysis (FMECA). This analysis led to the following main critical failure modes:

### v3

@@ Abstract
Artificial Intelligence , Attitude Control , CubeSats , Deep Reinforcement Learning , In-Orbit Demonstration , Nanosatellites
@@ 1 Introduction and Background
In recent years, Machine Learning (ML) and Artificial Intelligence (AI)-based methods have increasingly entered engineering disciplines, particularly robotics, with notable progress in quadrupedal [21] and humanoid robots such as Agile Justin developed at the German Aerospace Center (DLR) [20]. This influence is now expanding into the aerospace sector, where AI-driven methods are being explored for highly nonlinear applications, including robust attitude control for launch vehicles [31, 30], combustion chamber regulation [9], UAV control [8, 13], and satellite attitude control [29, 15, 19].
The motivation stems from the increasing autonomy and complexity required by modern spacecraft, which demand control strategies capable of adapting to dynamic environments while minimizing human intervention. Deep Reinforcement Learning (DRL), a variant of classical Reinforcement Learning [18], has shown considerable promise in this context. For example, for launch vehicles, DRL-based controllers using the Soft Actor–Critic [7] algorithm have been applied in simulation to achieve load-relief attitude control under aerodynamic disturbances [31]. Similarly, DRL-based adaptive controllers for quad-tiltrotor UAVs improve disturbance rejection and tracking performance compared to classical methods [8] in simulation.
For satellites, DRL has been used in simulation to maintain operational capability under unknown ADCS failures, where a Deep Deterministic Policy Gradient (DDPG) [12]-based controller reorients solar panels toward the Sun without requiring explicit fault models [29]. Other work demonstrates the feasibility of real-time DRL-based magnetic attitude control on embedded hardware on ground [15]. In addition, Proximal Policy Optimization (PPO) [24]-based controllers using stacked observations, similar to our observation space design from [10], achieve robust attitude control across varying satellite masses in simulation, outperforming classical PID control in settling time and stability [19]. The Hybrid Online Policy Adaptation Strategy (HOPAS) project attempted an early in-orbit demonstration of a Reinforcement Learning-based attitude controller on the ESA OPS-SAT CubeSat, with the AI agent designed to modify the PID control signals as a residual command, but the full AI closed-loop control was ultimately restricted to simulation and ground-testing [3, 1]. These developments illustrate the growing interest in DRL and AI-based control methods for aerospace systems characterized by nonlinear dynamics and uncertainties.
@@
Building on our previous works [10, 17, 11], we developed an AI-based attitude controller for the InnoCube 3U nanosatellite [2, 23, 27], developed by the Julius-Maximilians-Universität Würzburg and the Technische Universität Berlin, which successfully launched on January 14, 2025. The AI agent autonomously interacts with a simulated environment to learn a control strategy that maximizes a given reward function. The agent was first trained on a basic control task to obtain a base-agent, which was subsequently post-trained to produce the final flight-agent. The controller was trained entirely on ground, uploaded to InnoCube, and tested in orbit. Since the present work was done as part of the “In-Orbit Demonstrator Learning Attitude Control” (German: In-Orbit-Demonstrator Lernende Lageregelung; LeLaR) project, we also refer to the AI agent as LeLaR. The InnoCube project is funded by the Federal Ministry for Economic Affairs and Energy (BMWE) under grant number (FKZ) 50RU2000.
A disadvantage of DRL lies in the tendency of neural networks to achieve good results in interpolation tasks but failure in extrapolation problems [28]. In that sense, the process of transferring an AI agent trained in simulation to the real target system is called Sim2Real transfer and poses a significant challenge when applying DRL to real-world tasks. For practical applications, it is necessary to vary the boundary conditions during training, e.g., via domain randomization, to avoid so-called Out-of-Distribution errors. This includes, in particular, the moments of inertia, which we have already successfully considered in [10, 17] and utilize in this work.
@@ 2 Methodology
As base-agent, we employ the post-trained agent from our prior work [5], using a split policy network, consisting of a reaction wheel and a magnetorquer subnetwork. It was originally designed to avoid prolonged RW operation in the interval of [350,350][-350,350] rpm to avoid the inaccuracies stemming from the RWs rate estimator when operated in this range. Initial tests in-orbit however revealed that the RW subnetwork alone performed sufficient well, as we will discuss in Section 3. Along with changed safety requirements introduced after deployment but prior to the first test, this observation prompted us to use the RW subnetwork of the agent from [5] as base and to fine tune it through post-training for standalone use. We refer to the post-trained agent as LeLaR flight-agent or flight-agent in short.
Since [5] is available only in German, we adapt and summarize the essential elements required to understand the current work and expand on important aspects. This includes the base-agent design and the sensor calibration methodology adapted from our prior work, covering both the physical onboard sensors and the simulation environment used for training of all agents. It is to be noted, that sensor calibration is not a requirement introduced by the DRL procedure but should generally be performed for any satellite mission. It is included here for completeness, as calibration affects both the satellite’s sensors and the quality of the simulation. We then highlight the modifications to the model and reward function, as well as the updated operational requirements and boundary conditions, applied to obtain the flight-agent. Finally, the Safety Cage architecture is described, encompassing the AI agent to ensure safe and reliable operation in case of a fault.
@@ 2.1 InnoCube Satellite
InnoCube is a 3U CubeSat mission designed as a technology demonstrator for several innovative technologies and has successfully been launched on January 14, 2025 into a 520km sun-synchronous orbit. InnoCube carries two main experiments: “Skip The Harness” (SKITH), a fully wireless data bus for intra-satellite communication and “Fibre-Reinforced Spacecraft Wall for Storing Energy” (Wall#E), an experimental carbon fiber reinforced plastic solid-state battery. Secondary payloads include amateur radio payloads and the “Experiment for Precise Orbit Determination” (EPISODE) experiment, a software defined GNSS receiver. InnoCube is a collaborative effort between the Julius-Maximilians-Universität Würzburg and the Technische Universität Berlin [27]. The InnoCube satellite is a 3U CubeSat with a mass of 4.2 kg. An overview of the spacecraft’s subsystems is provided in Figure 1.
@@
On-board data handling is distributed across the satellite’s individual hardware nodes. The system is based on the Silicon Labs EFR32FG12 microcontroller, which features an ARM Cortex-M4 processor [25]. It integrates a 2.4 GHz RF front end that serves as the foundation for the wireless bus system Skip-the-Harness (SKITH). All spacecraft modules share this wireless data interface. Mission-critical bus functions are implemented within the PCDU, the COM node and the OBC node. These include power management, telemetry and telecommand functionality, as well as general bus operations.
@@ 2.1.1 InnoCube ADCS
As previously mentioned in Section 2.1, all spacecraft modules share the same EFR32FG12 microcontroller. On the ADCS mainboard, the basic circuitry of the microcontroller is extended to accommodate the electrical components needed for the ADCS. These include the gyroscope, the magnetometer, controllers and electrical interfaces for the magentorquers, electrical interfaces to the reaction wheels and other nescessary electrical systems. The used gyroscope to measure the body rates of the satellite is a STMicroelectronics ASM330LHH micro-electromechanical systems (MEMS) gyro [26]. The used magnetometer to measure the magnetic vector with respect to the body frame is a three-axis PNI RM3100 magneto-inductive sensor [16]. The external sun sensors are not used for the current LeLaR experiments, since we currently only determine the relative attitude using a quaternion integrator based on the measured gyro body rates.
@@ 2.2 Simulation Model
The simulated orbit is based on InnoCube TLE data from CelesTrak [4], with variations applied during training. Table 3 summarizes nominal values and offsets.
@@ 2.4.1 Sensor Noise Model
Bias std $\sigma_{b}$ (/s) $2.008\times 10^{-2}$ $1.419\times 10^{-2}$ $9.774\times 10^{-2}$
@@ 2.4.3 Residual Dipole Moment Estimation
Residual dipole $\mu$ and gyro bias $b$ were estimated from telemetry by solving a calibration problem based on Euler’s rotational dynamics [22]:
@@
where $\omega_{corrected}=\omega_{measured}+b$. The method was validated using simulated data with known ground truth and then applied to telemetry without actuator influence. Leveraging PyTorch [14] autograd gradient descent and the AdamW optimizer, we obtained:
@@ 2.5 Base-Agent Training
Using our AI agent from [5] as a base, we derive the flight-agent through post-training, which is conceptually related to transfer learning. The base-agent is initially trained to solve the general attitude control problem. Subsequently, it undergoes post-training with modified boundary conditions and a revised optimization target to produce the flight-agent. This approach leverages the well-initialized policy that already captures the fundamental dynamics of the system, avoiding the need to re-learn this knowledge as compared to training the flight-agent from scratch. As outlined in Section 1 the flight-agent is based off of the RW subnetwork of the base-agent. In the following section, the base-agent training is described. The modifications performed to obtain the flight-agent are explained in Section 2.6.
@@ 2.5.1 Base-Agent Reward Functions
The base-agent was trained with a PPO derivative, SkipPPO [5], using separate subnetworks for RWs and MTs. The RWs learn to quickly attain the goal attitude, while the MTs learn slow RW momentum management to bring the RW speeds to $\pm$500 rpm (corresponding to $\pm$350 rpm and a safety buffer). The RW reward function of the base-agent rewards quick reduction of attitude error, reduction of angular rates, and stable convergence. The MT reward encourages minimizing RW speed deviations. A combined reward was used during post-training of the base-agent to balance both effects. The RW reward function is shown in Equations (11)–(16).
@@ 2.7.1 Safety Cage
In order to assure that mission critical states resulting from unwanted controller behaviour are not reached during attitude control maneuvers, a Safety Cage architecture as proposed in [6] is implemented. This construct, recommended by the ECSS Machine Learning Handbook, uses a set of rules, to prevent the system from entering potentially dangerous states. This can be achieved through a set of safety interventions, which will be activated once a certain rule is triggered. Potentially critical failure modes for the InnoCube satellite have been identified using the proposed Failure Mode, Effects and Criticality analysis (FMECA). This analysis led to the following main critical failure modes: